Find the equation of the line through (4,7) Perpendicular to 2x+5y=15. answer And slope intercept
Answer by lwsshak3(11628)

Find the equation of the line through (4,7) Perpendicular to 2x+5y=15. answer And slope intercept
..
Equation of a line: y=mx+b, m=slope, b=y-intercept
Equation of given line:
2x+5y=15
5y=-2x+15
y=-2x/5+3
slope,m=-2/5
slope of line perpendicular to given line=5/2 (negative reciprocal)
equation: y=5x/2+b
solve for b using coordinates of given point (4,7)
7=5*4/2+b
7=10+b
b=-3
Equation of perpendicular line: y=5x/2-3